OpenStudy (anonymous):

Help Linear Equations Is (1, 5) a solution of y = 3x + 2?

Parth (parthkohli):

Substitute x with 1 and y with 5. Check for the answer then.

OpenStudy (anonymous):

Im thinking It is a solution but idk...

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow y = 3x + 2 \Longrightarrow 5 = 3(1) + 2 \Longrightarrow 5 = 3 + 2 \Longrightarrow5 = 5 \checkmark }\)

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

^Thanks

Parth (parthkohli):

yw lol

OpenStudy (anonymous):

lol :)